OpenStudy (anonymous):

What must be subtracted from 4x^4 + 2X^3 - 8X^2 + 3X - 7 so that it may be exactly divisible by 2x^2 + x - 2 ?

hero (hero):

First multiply 2x^2 + x - 2 by 2x^2: 2x^2(2x^2 + x - 2)

OpenStudy (mathstudent55):

Use polynomial long division to divide the polynomials. The answer is the remainder.

OpenStudy (anonymous):

thanku

OpenStudy (mathstudent55):

yw

hero (hero):

2x^2(2x^2 + x - 2) = 4x^4 + 2x^3 - 4x^2 4x^4 + 2X^3 - 8X^2 + 3X - 7 -4x^4 - 2x^3 + 4x^2 You subtract this: -4x^2 + 3x - 7 Because if you do that: 4x^4 + 2x^3 - 8x^2 + 3x - 7 - (-4x^2 + 3x - 7) = 4x^4 + 2x^3 - 8x^2 + 4x^2 + 3x - 3x + 7 - 7 = 4x^4 + 2x^3 - 4x^2 Which is divisible by 2x^2 + x - 2

OpenStudy (mathstudent55):

The way this problem is worded, there is not a unique solution. Let me explain what I mean with numbers, which we are more familiar with. We know that 57 is not divisible by 8. Let's say I ask the question "what must be subtracted from 57 to make it divisible by 8?" My approach was: Divide 57 by 8 to get a remainder of 1. Then subtract 1 from 57 to get 56 which is divisible by 8. My answer, then is 1, the remainder of the division. Another approach is: Multiply 8 by 2 to get 16. 16 is definitely a multiple of 8 and obviously divisible by 8. Then since 57 - 16 = 41, if you subtract 41 from 57 you get 16 which is divisible by 8. This was @Hero's approach. @Hero and I have two correct answers to the problem. By multiplying 2x^2 + x - 2 by 2x^2, Hero came up with a multiple of 2x^2 + x - 2 which is obviously divisible by 2x^2 + x - 2 because that is what a multiple is. My approach was to perform polynomial long division and come up with a remainder. Then if the remainder is subtracted from the dividend, then resulting polynomial is divisible by the divisor.